A company sells a phone for 120$ each for 10 or less. For each phone in excess of 10, the price is decreased 4$ per phone for everyone. What number of phone will produce the greatest revenue for the company?
Answer by Boreal(15235)

for x=1
that would be 116
for x=2 that would be 112
so revenue would be (10+x)(120-4x)
that is 1200+80x-4x^2 and has to be maximized
with x=-b/2a for the vertex, x is -80/-8=10
20 phones is the greatest revenue with 20*(120-40)=$1600
19 wold be 19*84=$1596
and 21 would be 21*76, also $1596